OpenStudy (anonymous):

help! i have been doing this for the past day.I need to pass this test. write the decimal 14.26848 as fraction.The 848 has a bar line on top,so it is a repeating decimal.It has to be in simplest form.If you can just write a mixed number instead of improper fractions. Thanks!

OpenStudy (anonymous):

help please!

OpenStudy (ranga):

Write it as: ___ x = 14.26848 First multiply both sides by powers of 10 such that the decimal point is one place to the RIGHT of the repeating digits. So if I multiply both sides by 100000 the equation becomes: 100000x = 1426848.(848 repeating) ----------- (1) Then multiply both sides so that the decimal point is 1 place to the LEFT of the repeating digits. Here, if I multiply by 100 the equation becomes: 100x = 1426.(848 repeating) ----------- (1) Subtract (2) from (1) and that will get rid of the repeating decimals. Then solve for x as a fraction.

OpenStudy (anonymous):

thank you so much.I will solve for x right now.

OpenStudy (ranga):

You are welcome. And after you get the fraction simplify it to the lowest fraction.

OpenStudy (anonymous):

okay.Thanks!

OpenStudy (ranga):

The problem says mixed numbers are okay and so I will just worry about putting the decimal into fraction form and leave 14 as it is.

OpenStudy (anonymous):

how do you solve for x?

OpenStudy (anonymous):

i don't know.

OpenStudy (ranga):

In the following, the digits within the parenthesis are indefinitely repeating digits. x = 0.26(848) 100000x = 26848.(848) 100x = 26.(848) subtract 99900x = 26822 divide both sides by 99900 x = 26822/99900 simplify. divide top and bottom by 2 x = 13411 / 49950 ( I don't think this fraction can be simplified any further So, 14.26(848) = 14 & 13411 / 49950 \[\Large 14.26(848) = 14\frac{ 13411 }{ 49950 }\] You can put just the fractional part in your calculator to verify that it gives ).26(848) The calculator will round off the last digit but that is fine.

OpenStudy (anonymous):

I asked my classmate how to do it.She gave me the same answer but didn't show me how to do it.She was mean to me.Thanks once again!

OpenStudy (ranga):

you are welcome.